# KACQ-DCNN: Uncertainty-Aware Interpretable Kolmogorov–Arnold Classical–Quantum Dual-Channel Neural Network for Heart Disease Detection

### v1

@@ Conclusions
In this study, we propose the KACQ-DCNN model, which is a novel quantum-classical hybrid architecture that combines KAN with CQML techniques. This model has demonstrated promising results in heart disease diagnosis. Our comprehensive statistical analysis revealed that the KACQ-DCNN 1-Layered model outperformed 34 benchmark models, including 16 CML models, 12 QNN models, and 6 hybrid models. Specifically, our KACQ-DCNN 1-Layered model achieved an accuracy of $92.03\%\pm 0.19\%$, a maP of $92.00\%\pm 0.19\%$, a maR of $92.00\%\pm 0.10\%$,and a maF1 score of $92.00\%\pm 0.07\%$. Additional metrics further support the model’s effectiveness, with a ROC-AUC score of $94.77\%\pm 0.43\%$, MCC of $83.85\%\pm 0.81\%$, and Cohen’s kappa of $83.84\%\pm 0.43\%$. The robustness of these results was confirmed via stratified 10-fold cross-validated two-tailed paired t-tests. Statistical significance was established via the Bonferroni correction, with an $\alpha_{\text{adjusted}}$ value of $0.0056$. All $p$-values were substantially lower than this threshold, allowing us to reject the $H_{0}$ for all models and confirming the superiority of KACQ-DCNN. Our ablation study highlights the synergistic effects of the hybrid approach. The ROC-AUC column shows a significant variance in the ablation test, but KACQ-DCNN 1-Layered outperforms all the other metrics when every component works in harmony.

### v3

@@ Conclusions
In this study, we propose the KACQ-DCNN model, a novel classical-quantum hybrid architecture that replaces MLP with KAN layers, demonstrating promising results in heart disease diagnosis. Our comprehensive benchmarking revealed that the KAN-enhanced KACQ-DCNN 1-Layered model outperformed 34 benchmark models, including 16 CML models, 12 QNN models, and 6 hybrid models. Specifically, our KACQ-DCNN 1-Layered model achieved an accuracy of $92.03\%\pm 0.19\%$, a maP of $92.00\%\pm 0.19\%$, a maR of $92.00\%\pm 0.10\%$, and a maF1 score of $92.00\%\pm 0.07\%$. Additional metrics further support the model’s effectiveness, with a ROC-AUC score of $94.77\%\pm 0.43\%$, MCC of $83.85\%\pm 0.81\%$, and Cohen’s kappa of $83.84\%\pm 0.43\%$. The robustness of these results was confirmed via stratified 10-fold cross-validated two-tailed paired t-tests. Statistical significance was established via the Bonferroni correction, with an $\alpha_{\text{adjusted}}$ value of $0.0056$. All $p$-values were substantially lower than this threshold, allowing us to reject the $H_{0}$ for all models and confirming the superiority of KACQ-DCNN. Our ablation study highlights the synergistic effects of the hybrid approach. The ROC-AUC column shows a significant variance in the ablation test, but the KAN-version of KACQ-DCNN 1-Layered outperforms all the other metrics when every component works in harmony. Although the dual-channel design choice may diminish the architecture’s effectiveness in real-time scenarios, the versatility of the KACQ-DCNN and its capacity to handle complex datasets with high precision makes it an optimal choice for our objectives.